An electrochemical cell is based on the following two half-reactions: Ox: Pb(s)→Pb2+(aq, 0.29 M )+2e− Red: MnO−4(aq, 1.30 M )+4H+(aq, 1.7 M )+3e−→ MnO2(s)+2H2O(l) show work
Answer #1

Consider reaction taking place in the cell.

At anode: Pb (s) Pb 2+ (aq) + 2 e-

At cathode : MnO 4-(aq) + 4 H + + 3 e- MnO2(s) + 2 H2O (l)

Overall reaction : Pb (s) + MnO 4-(aq) + 4 H + + 3 e- Pb 2+ (aq) + 2 e- +  MnO2(s) + 2 H2O (l)

Balance reaction for electrons, we get

3 Pb (s) +2 MnO 4-(aq) + 8 H + 3 Pb 2+ (aq) +2 MnO2(s) + 4 H2O (l)

Here no of electrons transferred = 6

Standard emf of the cell is calculated as

E 0 cell = E 0 cathode - E 0 anode  

= 1.692 - ( -0.126 )

= 1.692+0.126

= 1.818 V

Emf of cell is calculated by using Nerns't equation as

E cell = E 0 cell - 2.303 R T / n F log [ Pb 2+ ] 3 / [MnO 4-  ] 2 [H + ] 8

We have 2.303 R T / F = 0.0591 at 25 0 C then

E cell = E 0 cell - 0.0591 / n log [ Pb 2+ ] 3 / [MnO 4-  ] 2 [H + ] 8

= 1.818 - 0.0591/ 6 log ( 0.29) 3 / (1.30) 2 ( 1.70) 8

= 1.818 - 0.00985 log 2.069 x 10 -04

= 1.818 - 0.00985 x ( - 3.684)

= 1.818 + 0.0363

= 1.854 V